OpenStudy (anonymous):

Will fan and medal. Question already answered, just need someone to check my work. :) Anyone?

OpenStudy (anonymous):

i will do it wat u need

OpenStudy (anonymous):

I got 83/100 but I'm not sure.

OpenStudy (anonymous):

whats the question

OpenStudy (igreen):

No..you have to divide.

OpenStudy (igreen):

Divide: \(\sf\dfrac{3}{4} \div 9\)

OpenStudy (anonymous):

yea divide

OpenStudy (igreen):

Or multiply by 9's reciprocal: \(\sf \dfrac{3}{4} \times \dfrac{1}{9} \rightarrow \dfrac{3 \times 1}{4 \times 9}\) Simplify

OpenStudy (anonymous):

Ok so I got 3/36

OpenStudy (anonymous):

yep

OpenStudy (igreen):

Yes, now simplify it.

OpenStudy (igreen):

\(\sf\dfrac{3}{36} \rightarrow \dfrac{3\div 3}{36 \div 3}\)

OpenStudy (anonymous):

0.083 i think she was right at the beginning

OpenStudy (anonymous):

yea she was

OpenStudy (igreen):

No..she said 83/100..that's 0.83

OpenStudy (igreen):

Plus, 3/36 is equal to \(\sf 0.08\overline 3\).

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

yea i kno

OpenStudy (anonymous):

Sorry I was gone for a minute. I got 1/12 when I simplified it

OpenStudy (anonymous):

that should be your answer

OpenStudy (anonymous):

Thank you :)

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

so who got the fan and medal